OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are given. A (2, 3) and B (6, 11)

OpenStudy (anonymous):

this is not too bad take the average in each coordinate i.e. add and divide by 2

OpenStudy (anonymous):

Would I use the distance formula to solve?

OpenStudy (anonymous):

no it seems like you would have to, because it involves distance, but it is actually much simpler

OpenStudy (anonymous):

okay! i'll calculate it

OpenStudy (anonymous):

what is the average of \(2\) and \(6\) ?

OpenStudy (anonymous):

would it be 4? (2+6)/2

OpenStudy (anonymous):

yes, the first coordinate of the midpoint is 4 how about the second coordinate?

OpenStudy (anonymous):

8.5

OpenStudy (anonymous):

no i don't think so try again, maybe you added wrong

OpenStudy (anonymous):

I keep getting 8.5

OpenStudy (anonymous):

(11+6)/2

OpenStudy (anonymous):

ooh i see what you did take the average of the second coordinates, not the average of 6 and 11

OpenStudy (anonymous):

\[(x_1,y_1), (x_2,y_2)\] the midpoint, compute the average of \(x_1\) and \(x_2\) i.e. \(\frac{x_1+x_2}{2}\) for the fist coordinate the second coordinate is \(\frac{y_1+y_2}{2}\)

OpenStudy (anonymous):

11

OpenStudy (anonymous):

\(\frac{3+11}{2}\) is ?

OpenStudy (anonymous):

\[ (2, 3) (6, 11)\] first coordinate of the midpoint is \[\frac{2+6}{2}\] and the second coordinate is \[\frac{3+11}{2}\]

OpenStudy (anonymous):

alrighty! thank you!:)

OpenStudy (anonymous):

yw